OpenStudy (anonymous):

Please help! Show all steps. Will give medal and fan to best answer! What is the graph of 4x^2+y^2=16? What are its lines of symmetry? What are the domain and range?

OpenStudy (anonymous):

it's an ellipse, draw a diagram and you'll see everything

OpenStudy (anonymous):

I don't understand how to do that. @whpalmer4

OpenStudy (rational):

divide the whole equation by 4

OpenStudy (anonymous):

x^2+y^2=4?

OpenStudy (rational):

sorry divide the whole equation by 16, so that right hand side becomes 1 : \(\large 4x^2+y^2=16\) \(\large \frac{4x^2}{16}+\frac{y^2}{16}=1\) \(\large \frac{x^2}{4}+\frac{y^2}{16}=1\) \(\large \frac{x^2}{2^2}+\frac{y^2}{4^2}=1\)

OpenStudy (rational):

thats a standard equation of ellipse centered at (0, 0)

OpenStudy (rational):